The difference between the measures of two supplementary angles is 102 degrees. Write an expression and find the measure if each angle.
Answer by Boreal(15235)

They are x and 180-x
180-x-x=102
-2x=-78
x=39 degrees
180-x=141 degrees. They are 102 degrees apart.